# The Tension of Space as Dark Energy: A No Geometric Sequestering Theorem for Minimal Sequesters

### v2

@@ 5 Discussion
We developed a minimal framework in which space is modeled as an elastic three membrane that fills spacetime. The dynamics are captured by an intrinsic Nambu–Goto volume term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$. This fixes the geometric stress tensor to $T^{(s)}_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ and identifies the membrane tension with a uniform vacuum energy of space. We then embedded this term into a matter vacuum sequester framework that employs global four form constraints to cancel constant contributions from the matter sector. The variation shows that the global variable adjusts to neutralize the matter vacuum energy while a purely geometric unit operator (the operator which behaves like vacuum energy) carries no dependence on that variable and is not in the cancellation channel. As a result the local field equations retain a residual contribution $T_{s}\,g_{\mu\nu}$. We interpret this as the effective cosmological term.
Subsequently we performed the path integral over the gravitational fluctuation $h_{\mu\nu}$ and over the associated Faddeev–Popov ghosts only. We did not integrate over matter fluctuations so as to not “recompute” the pure matter loops. Their effect is assumed to have already been absorbed into the renormalized matter functional $S_{m}[\lambda^{-2}g,\psi]$ at the matching scale $\mu$.
The outcome was a background diffeomorphism invariant one particle irreducible action $\Gamma[\bar{g}]$ with the structure shown in equation (4.15). The local part is an infinite covariant tower built from the curvature and its derivatives. The non-local part contains form factors that never reduce to a pure volume term and so do not influence vacuum energy. The geometric volume tension $T_{s}(\mu)$ arises from the renormalized coefficient of the unit operator and enters the field equations. This separation makes clear which ingredients can shift $T_{s}(\mu)$ and which ingredients only change higher derivative responses of the geometry.
If one wishes to include matter-graviton mixed loops explicitly, one can extend the path integral to the matter fields fluctuations as well. The background field method and general covariance then enforce the same covariant operator basis that appears in equation (4.15). Mixed matter-graviton diagrams that contain at least one internal graviton line are local in the ultraviolet. They therefore renormalize the same coefficients that multiply curvature invariants as in equation (4.15). No new local scalar of lower canonical mass dimension exists beyond the unit operator and the Ricci scalar term. As a result the only change is in the numerical values of the Wilson coefficients.
Concretely, mixed matter–graviton loops shift the same $c_{0}(\mu)$ and therefore in turn shift $T_{s}(\mu)\equiv T_{s}^{\rm bare}-c_{0}(\mu)$. They also shift $c_{2}(\mu)$ which renormalizes the constant $G$ along with shifting the coefficients of curvature squared invariants and beyond. The tensorial form of each contribution is fixed by background diffeomorphisms and does not depend on the details of the gauge fixing. It can further be arranged so that pure matter functional $S_{m}[\lambda^{-2}g,\psi]$ will still continue to encode matter self interactions and pure matter loops at the chosen matching scale. The combined result is a single covariant effective action with renormalized coefficients $c_{0}(\mu),c_{1}(\mu),\cdots$ that capture both gravitational and mixed contributions in a uniform way. Therefore, the generalization to mixed matter-graviton graphs is straightfoward.
With this in mind, one can then conveniently regard $T_{s}(\mu)$ as the fully renormalized geometric tension that already includes contributions from pure graviton self loops and from mixed matter–graviton loops because as mentioned before, extending the path integral from metric fluctuations to metric plus matter fluctuations only shifts the Wilson coefficients in the same covariant operator basis. The definition of $T_{s}(\mu)$ as the renormalized coefficient of the unit operator therefore captures all such shifts in a single parameter.
One might try to redefine $\Lambda^{\prime}\equiv\Lambda-T_{s}$ so that the sequestering constraint cancels both the matter vacuum energy and the geometric volume term. This identification is physically not possible because in the standard sequestering framework which we are considering the global variables and four-forms are wired so that the variational constraint removes constant shifts of the matter Lagrangian, while the purely gravitational sector is inert under that constraint. An independent geometric unit operator $T_{s}\int\sqrt{-g}$ is therefore not neutralized unless one explicitly extends the mechanism to couple the same global sector to that operator. To also cancel $T_{s}$, one would need a different theory, one which couples a new rigid variable to a geometric density. This represents a different sequestering scheme and whereas we are working in the minimal framework which requires just two global Lagrange multipliers $\lambda$ and $\Lambda$. One can also see why such an identification is not possible in the minimal case because the operator $T_{s}\int\sqrt{-g}$ is blind to $\lambda$ and therefore cannot contribute to the $\lambda$ constrain.
So it is correct to keep $\Lambda$ as the variable that enforces the global matter constraint and we keep $T_{s}(\mu)$ as an intrinsic geometric coupling that lives in the gravitational effective action. In the minimal sequester framework, the added advantage of having an un-sequestered geometric term is that after fixing covariant operator basis using the background field method, the term $T_{s}(\mu)$ automatically emerges as the renormalized coefficient of the unit operator after graviton and mixed matter–graviton loops are included. Nonlocal form factors appear but they do not produce a pure volume term and therefore do not modify this identification. The constant source in the field equations can therefore naturally be identified as $\rho_{\rm cav}(\mu)=T_{s}(\mu)$. At an infrared scale $\mu_{\rm IR}$ appropriate for cosmology one can then match $\rho_{\rm vac}(\mu_{\rm IR})$ to the observed dark energy density. This retains the virtue of sequestering for matter vacuum energy while leaving a natural geometric volume term that can account for the residual dark energy. It also avoids promoting geometric couplings to global multipliers and keeps conservation and locality manifest through the Bianchi identity.
@@ 6 Conclusion
We have modeled space as an elastic three membrane and identified its uniform volume tension with a geometric unit operator that contributes a stress proportional to the metric. The Nambu–Goto term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ fixes $T^{(s)}_{\mu\nu}=-T_{s}g_{\mu\nu}$ and allows a direct comparison with vacuum energy. Embedding this structure into a four–form sequester framework, we showed that global constraints neutralize constant matter contributions inside the invariant combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$, while the geometric unit operator does not couple to the global variable and therefore cannot be removed. The local field equations then take the renormalized effective form,
$M_{*}^{2}(\mu)\,G_{\mu\nu}=\bigl[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\bigr]-T_{s}(\mu)\,g_{\mu\nu}+H_{\mu\nu}[R^{2},\nabla R,\ldots]$
with higher–derivative tensors suppressed at low curvature. Using the background field method we argued that graviton loops and mixed matter–graviton loops renormalize the unit operator, so $T_{s}(\mu)$ runs with the scale. The residual constant source may be identified with dark energy at an infrared matching scale relevant for cosmology.
An immediate next step is to extend the elastic space picture by introducing bulk modulus and bulk viscosity to space. In a homogeneous expansion a nonzero bulk viscosity produces an effective negative pressure that can evolve with the Hubble rate. This offers a natural route to time dependent dark energy and can help account for DESI indications of evolving dark energy. We will develop this viscous elastic space framework next.
@@ 7 Acknowledgment
The author acknowledges the use of ChatGPT-5 (August 2025 version) to edit and refine tone and clarity in selected sections of this manuscript. The model was used only to polish and rephrase text supplied by the author. All LLM assisted passages were reviewed and edited by the author, exercising the final editorial judgment.

### v3

@@ 5 Discussion
The present construction should be viewed as a proof of concept for a particular physical idea, namely that the late-time dark-energy sector may be interpreted as the intrinsic tension of space itself, and that mild departures from exact vacuum behavior may arise if a hidden defect reservoir exchanges energy with the dynamical part of that tension. In this framework, the observed cosmological term is not replaced by an ad hoc dark-energy fluid. Rather, one begins from an intrinsic membrane description of space, identifies its uniform tension with the residual vacuum sector, and then asks whether additional hidden-sector structure can perturb that vacuum contribution away from exact constancy at late times.
A central outcome of the analysis is that the model does generate a nontrivial effective equation of state for dark energy. In particular, once the total tension is split into a rigid background piece and a dynamical correction, the exchange of energy between the hidden flux-tube reservoir and the latter induces a time-dependent $w_{\rm eff}(a)$. At the coarse-grained level, this setup also permits a transient crossing of the phantom divide. In this sense, the model succeeds in establishing the basic logical point that a running tension of space can mimic a mild low-redshift evolution of the dark-energy sector without requiring the vacuum baseline itself to be removed or destabilized.
The benchmark fit discussed in Section 4 also offers a useful piece of physical guidance. Within the restricted family considered there, the preferred solution is driven toward the earliest transition epoch allowed by the DESI leverage range and toward the largest possible initial hidden flux fraction. Moreover, the best restricted fit corresponds to $n=2$, which implies $w_{\rm flux}=-1/3$. This is not radiation-like behavior. Rather, it is the equation of state expected of an effectively string-like network of line-like defects. That result is physically consistent with the picture developed in Section 3, where the late-time broken phase traps hidden magnetic flux into tube-like configurations instead of leaving it in a bath of freely propagating relativistic quanta. The fit therefore points back toward the underlying interpretation of the model, namely that a sufficiently persistent defect reservoir is favored if one wishes to generate an observable running of the tension sector.
At the same time, the present treatment has important limitations. The first and most obvious one is that the hidden flux-tube network has not been modeled in detail. The late-time $U(1)_{h}\to\mathbb{Z}_{n}$ transition is introduced only at the effective level, and the subsequent defect formation and full network evolution are not computed microscopically. Instead, all of that complicated nonequilibrium dynamics is compressed into the effective quantities $\rho_{\rm flux}$ and $w_{\rm flux}$. This simplification is appropriate for a first proof-of-concept study, but it also means that the cosmological sector of the model is not yet tied to a fully derived defect-network history.
A second limitation is that the exchange term between the hidden reservoir and the dynamical part of the tension sector is phenomenological. The form adopted in this work is simple, physically interpretable, and sufficient to illustrate the desired effect, but it has not been derived from a first-principles treatment of the local hidden-sector action. A more complete analysis would need to show how the effective transfer law emerges from the microscopic dynamics of the hidden Higgs field, the hidden gauge sector, and the evolving defect network. Closely related to this is the fact that the late-time phase transition itself is assumed rather than derived from a detailed cosmological history. In particular, the present work does not yet explain why the hidden-sector transition should occur precisely in the redshift interval where a DESI-like signal would be most relevant.
A third limitation concerns the observational comparison. The analysis of Section 4 does not fit the full BAO, supernova, and CMB likelihoods directly. Instead, it compares the model against a compressed Gaussian description of the effective CPL posterior in the $(w_{0},w_{a})$ plane. This makes the comparison transparent and reproducible, and it is entirely adequate for a first diagnostic test of the idea. However, it also means that the resulting agreement should not be interpreted as a precision cosmological fit. The residual gap between the best-fit model curve and the benchmark target is therefore not surprising. Indeed, that mismatch should be read as a sign that the minimal single-reservoir ansatz is still too restrictive, not as evidence that the underlying physical idea must fail. There is substantial room for improvement once the hidden-sector dynamics, the network evolution, and the exchange mechanism are treated in greater detail.
There is also an important conceptual limitation in scope. The present model does not solve the cosmological constant problem itself. The quantity $T_{s}$ is interpreted throughout as the small residual vacuum sector that remains after whatever ultraviolet physics is responsible for suppressing the enormous microscopic vacuum contributions has already done its job. In other words, the present framework is not a first-principles derivation of why the vacuum energy is tiny. Rather, it is an attempt to understand whether the surviving late-time vacuum sector may admit a natural effective description in terms of an intrinsic space tension, and whether hidden-sector dynamics can generate a mild departure from exact constancy around that already-small baseline.
Finally, the present work is restricted to the homogeneous background evolution. It does not yet address perturbations in the hidden sector, the clustering properties of the defect reservoir, possible stability issues beyond the background level, or additional signatures that may arise outside the effective CPL plane. These questions are essential if the framework is to be developed into a more complete phenomenological model.
Taken together, these limitations define a clear program for future work. The most immediate next step is to replace the coarse-grained reservoir description by a more explicit treatment of the hidden defect network and its late-time cosmological evolution. A second step would be to derive, or at least better motivate, the exchange channel from the underlying local action. A third would be to extend the analysis from the compressed CPL benchmark to a full likelihood-based confrontation with cosmological data. It is also worth noting that the tension-of-space framework appears flexible enough to support more than one late-time dynamical realization. In particular, the companion phonon model of Ref. [55] implements the same broad geometric viewpoint through a viscous longitudinal phonon sector and was found to closely track a DESI-motivated CPL benchmark over the DESI-sensitive redshift range.
@@ 6 Conclusion
In this work, we have explored the possibility that the late-time dark-energy sector may be interpreted as the intrinsic tension of space itself. Starting from an intrinsic membrane description, we showed that a uniform space tension contributes to the gravitational field equations with exactly the tensor structure of vacuum energy. We then considered the Dirac–Born–Infeld completion of this picture, which naturally introduces a hidden Abelian gauge sector in addition to the vacuum-like tension term.
[…]
We then compared the resulting evolution with a compressed observational benchmark in the effective CPL plane. Although the minimal single-reservoir realization does not exactly reproduce the benchmark target, it is able to generate a nontrivial late-time running that lies in the phenomenologically relevant direction. In particular, the preferred restricted fit points toward an early onset of the transition within the DESI-sensitive range and toward a substantial hidden defect population, which is consistent with the basic physical picture developed in the model.
The present framework should be interpreted as a proof of concept rather than as a complete theory of dark energy. We have not attempted to derive the late-time hidden-sector transition from first principles, nor have we modeled the full formation and evolution of the hidden flux-tube network. Likewise, the exchange law between the defect reservoir and the tension sector has been treated phenomenologically, and the observational comparison has been carried out at the level of a compressed CPL benchmark rather than through a full likelihood analysis. Even so, the construction demonstrates a logically consistent and physically suggestive possibility, in that once the vacuum sector is reinterpreted as an intrinsic tension of space, hidden-sector defect physics can naturally induce a mild running of the effective dark-energy component.
There are several natural directions for future work. A more complete treatment should derive the effective transfer mechanism from the underlying hidden-sector dynamics, model the defect-network evolution in greater detail, and extend the present background-level study to perturbations and full cosmological likelihood tests. It would also be interesting to investigate whether modest extensions of the present setup can reduce the remaining gap to the observational benchmark. For these reasons, we regard the present model as a minimal starting point for a broader program in which the dark-energy sector is viewed not as an independent fluid, but as a dynamical manifestation of the intrinsic tension of space.
@@ 7 Acknowledgment
The author acknowledges the use of ChatGPT-5.4 to edit and refine tone and clarity in selected sections of this manuscript. The model was used only to polish and rephrase text supplied by the author. All LLM assisted passages were reviewed and edited by the author, exercising the final editorial judgment.
Appendix A Derivation of the effective equation of state
We now derive the relation,
$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln T}{d\ln a}\,.$ (A.1)
The starting point is the continuity equation for a homogeneous effective fluid in an FLRW background,
$\dot{\rho}_{\rm DE}+3H\bigl(\rho_{\rm DE}+p_{\rm DE}\bigr)=0\,$, (A.2)
where $H=\dot{a}/a$ is the Hubble parameter. Writing the pressure in the usual form,
$p_{\rm DE}=w_{\rm eff}\,\rho_{\rm DE}\,$, (A.3)
one finds,
$\dot{\rho}_{\rm DE}+3H\bigl(1+w_{\rm eff}\bigr)\,\rho_{\rm DE}=0\,.$ (A.4)
Solving for $w_{\rm eff}$ gives,
$w_{\rm eff}=-1-\frac{\dot{\rho}_{\rm DE}}{3H\rho_{\rm DE}}\,.$ (A.5)
Using,
$\frac{d\ln\rho_{\rm DE}}{d\ln a}=\frac{1}{\rho_{\rm DE}}\,\frac{d\rho_{\rm DE}}{d\ln a}=\frac{\dot{\rho}_{\rm DE}}{H\rho_{\rm DE}}\,$, (A.6)
Equation (A.5) therefore becomes,
$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln\rho_{\rm DE}}{d\ln a}.$ (A.7)
In the present framework, the effective dark-energy density is identified with the tension of space,
$\rho_{\rm DE}(a)\equiv T(a)\,.$ (A.8)
Therefore,
$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln T}{d\ln a}\,.$ (A.9)
If we further split the tension as,
$T(a)=T_{s}+\delta T_{s}(a)\,,\qquad\dot{T}_{s}=0\,$, (A.10)
then,
$\frac{d\ln T}{d\ln a}=\frac{1}{T_{s}+\delta T_{s}}\frac{d\delta T_{s}}{d\ln a}\,.$ (A.11)
Substituting this into Eq. (A.9), one obtains,
$w_{\rm eff}(a)=-1-\frac{1}{3\bigl(T_{s}+\delta T_{s}\bigr)}\frac{d\delta T_{s}}{d\ln a}\,.$ (A.12)
Appendix B Derivation of the coupled continuity equations
We model the late-time interacting sector in coarse-grained form as the sum of two components, fist of which is a hidden flux-tube reservoir and second being the dynamical part of the space-tension sector. Accordingly, we split the total stress tensor as,
$T^{\mu\nu}_{\rm tot}=T^{\mu\nu}_{\rm flux}+T^{\mu\nu}_{\delta T_{s}}\,.$ (B.1)
We assume that the total interacting sector is covariantly conserved,
$\nabla_{\mu}T^{\mu\nu}_{\rm tot}=0\,$, (B.2)
but that the two individual components may exchange energy through an interaction current $Q^{\nu}$ according to,
$\nabla_{\mu}T^{\mu\nu}_{\rm flux}=-Q^{\nu}\,,\qquad\nabla_{\mu}T^{\mu\nu}_{\delta T_{s}}=+Q^{\nu}\,.$ (B.3)
For a homogeneous FLRW background, the transfer is taken to be purely temporal in the comoving frame,
$Q^{\nu}=Q\,u^{\nu}\,$, (B.4)
where $u^{\nu}$ is the background four-velocity and $Q$ is the scalar energy-transfer rate.
The flux-tube sector is described as an effective fluid with,
$p_{\rm flux}=w_{\rm flux}\,\rho_{\rm flux}\,.$ (B.5)
Its continuity equation then becomes,
$\dot{\rho}_{\rm flux}+3H\bigl(\rho_{\rm flux}+p_{\rm flux}\bigr)=-\,Q\,$, (B.6)
or equivalently,
$\dot{\rho}_{\rm flux}+3H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}=-\,Q\,.$ (B.7)
We now treat the dynamical transferred part of the space tension as a vacuum-like component with,
$\rho_{\delta T_{s}}=\delta T_{s}\,,\qquad p_{\delta T_{s}}=-\,\delta T_{s}\,.$ (B.8)
Its continuity equation is therefore,
$\dot{\rho}_{\delta T_{s}}+3H\bigl(\rho_{\delta T_{s}}+p_{\delta T_{s}}\bigr)=Q\,.$ (B.9)
Since,
$\rho_{\delta T_{s}}+p_{\delta T_{s}}=0\,$, (B.10)
this reduces immediately to,
$\dot{\delta T_{s}}=Q\,.$ (B.11)
We thus obtain the coupled background system,
$\dot{\rho}_{\rm flux}+3H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}$ $=-\,Q\,$, (B.12)
$\dot{\delta T_{s}}$ $=Q\,.$ (B.13)
These equations simply express that the hidden flux-tube sector loses exactly the amount of energy gained by the dynamical part of the space-tension sector, and conversely when the sign of $Q$ reverses.